# Synthetic vs. Gold: The Role of LLM-Generated Labels and Data in Cyberbullying Detection

### v2

@@ Abstract
Cyberbullying (CB) presents a pressing threat, especially to children, underscoring the urgent need for robust detection systems to ensure online safety. However, progress in developing such systems is hindered by the scarcity of large, labeled datasets that are specifically tailored for specialized tasks and the target age groups. Creating these datasets relies heavily on human annotation, which not only strains resources but also raises significant ethical and legal concerns due to annotators’ exposure to harmful content, not withstanding the acquisition of this type of data from vulnerable populations such as children. In this paper, we address these challenges by leveraging Large Language Models (LLMs) to generate synthetic data and labels. Our experiments demonstrate that synthetic data enables BERT-based CB classifiers to achieve performance close to that of those trained on fully authentic datasets (75.8% vs. 81.5% accuracy). Additionally, LLMs can effectively label authentic yet unlabeled data, allowing BERT classifiers to attain a comparable performance level (79.1% vs. 81.5% accuracy). These results highlight the potential of LLMs as a scalable, ethical, and cost-effective solution for generating data for CB detection.
Synthetic vs. Gold: The Role of LLM-Generated Labels and Data in Cyberbullying Detection
Arefeh Kazemi and Sri Balaaji Natarajan Kalaivendan and Joachim Wagner and Hamza Qadeer and Brian Davis School of Computing, ADAPT Centre, Dublin City University, Dublin, Ireland Correspondence: firstname.lastname@adaptcentre.ie
@@ 1 Introduction
The rapid proliferation of social media platforms has raised concerns over the prevalence of cyberbullying (CB), particularly among vulnerable populations like children. Detecting and mitigating CB is crucial to maintaining a safe digital environment and minimizing its psychological impact Arif et al. (2024). However, creating high-quality labeled datasets for training CB detection models remains a significant challenge. Traditional data collection methods, which are based on human annotators, are costly, time-consuming, and pose ethical and legal concerns. Annotators may experience emotional distress or harm when exposed to harmful content, raising questions about their well-being AlEmadi and Zaghouani (2024). This highlights the need for alternative approaches to generate labeled datasets without requiring human annotators to engage directly with harmful content. One promising solution is to use large language models (LLMs) for the annotation of data, instructing an instruct-tuned model to perform the task.
Since detecting CB is particularly important in protecting children online, our industry partner aims to integrate built-in child-protection software into children’s smartphones to ensure safer internet use. By analyzing incoming and outgoing data on a child’s device, the software filters offensive content, including CB, and alerts parents if any imminent danger is detected. As part of this initiative, we investigate the use of LLMs to create datasets for CB detection. However, collecting CB content produced by children is difficult. Age information on public social media sites is unreliable and a large part of CB happens in closed groups. Furthermore, obtaining consent involves the parents, who may not be interested in (or may not have time to consider) taking part in a research study. A solution may be to use LLMs to generate conversations between children.
LLMs have shown remarkable capabilities to generate human-like text and perform various NLP tasks, including text classification and sentiment analysis Brown et al. (2020). These models can generate synthetic data that mimic authentic content and, as automatic annotators, alleviate both the financial and ethical burdens associated with manual annotation. Recent studies have demonstrated the utility of LLM-generated data in NLP tasks where authentic datasets are scarce or difficult to acquire Li et al. (2023); He et al. (2022); Ghanadian et al. (2024); Yoo et al. (2021).
@@ 2 Background
In recent years, the use of LLMs for synthetic data generation has gained significant traction due to its efficiency, scalability, and cost-effectiveness compared to human-annotated datasets. He et al. (2021, 2022) generate synthetic data for knowledge distillation, self-training, and few-shot learning tasks, which is then annotated using state-of-the-art classifiers. Bonifacio et al. (2022) employ LLMs to generate labeled data in a few-shot manner for information retrieval. Yoo et al. (2021) generate augmented text samples by selecting a few sentences from task-specific training data and embedding them into prompts. Anaby-Tavor et al. (2020) fine-tune a pre-trained language model on a small labeled dataset. The fine-tuned model is then used to generate new labeled text samples. Meng et al. (2022) employ an LLM to generate class-conditioned texts based on label-descriptive prompts for a classification task.
While the aforementioned studies focus on general NLP tasks, some research explores synthetic data generation for specific domains. harmful-content detection. Wang et al. (2024) introduce NoteChat, a multi-agent framework designed to generate synthetic patient-physician conversations from clinical documents. Ghanadian et al. (2024) create socially aware synthetic datasets for suicidal ideation detection. Previous studies exploring the use of synthetic data have reported mixed results on whether LLM-generated synthetic data can effectively train models to perform at a level comparable to those trained on authentic data Li et al. (2023).
There also is research on generating semi-synthetic data for CB detection. Ejaz et al. (2024) generate a semi-synthetic CB dataset by creating synthetic users. Aggressive and non-aggressive messages are sourced from existing authentic datasets and randomly assigned to user interactions. CB instances are labelled using a threshold-based approach, flagging interactions as CB if they meet predefined criteria for peerness, intent to harm and message repetition. This study creates a dataset, trying to simulate real CB behaviors, where the users are semi-synthetic, and the messages are generated using a combination of existing authentic CB datasets and synthetic data. Pérez et al. (2024) generate a synthetic CB dataset using a Bayesian Network (BN) model Pearl (1988) trained on real-world survey data collected from minors. To generate synthetic data, the Bayesian Model Sampling algorithm was applied to the trained BN, producing risk profiles for synthetic agents. The dataset consists of synthetic players with demographic attributes and responses to 15 binary questions related to cyber risk.
@@
In contrast to these works, our study aims to explore the use of LLMs to generate a fully synthetic CB dataset or to label existing authentic but unlabeled CB datasets. Our work uniquely applies LLM-generated data to CB detection, an area where data collection is both ethically and logistically challenging. By assessing the effectiveness of LLM-generated datasets in comparison to authentic data, we provide new insights into the viability of LLMs in training models for sensitive tasks like CB detection.
@@ 3.1 Overview of Scenarios
We investigate the role of LLMs in CB detection, focusing on their utility under varying data availability conditions and under the assumption that direct use of LLMs as a classifier is too expensive due to the high volume of messages to be checked. As a baseline for comparison, we evaluate a scenario in which a lightweight, BERT-based classifier is trained exclusively on gold-standard, manually labeled authentic data without LLM involvement. We then define three additional scenarios with different data availability and that use LLMs in different ways.
@@
This scenario represents the ideal situation where sufficient gold-standard data is available for fine-tuning BERT. It serves as the benchmark for evaluating the effectiveness of other approaches. In this setup, no synthetic data or LLMs are involved. This scenario is feasible if resources such as time, budget and expert annotators are abundant. However, it often proves impractical due to the challenges of manual labeling.
@@
This scenario applies when labeled authentic data is unavailable, and there is no intention to train a separate classifier for CB detection. Instead, an instruction-tuned LLM is used directly as a classifier, leveraging its pre-trained knowledge and its ability to follow instructions to identify CB instances. The primary advantage of this method is its elimination of the need for labeled data and training time. However, there are trade-offs. While an LLM can handle nuanced language patterns, it may be less efficient and incur higher computational costs compared to simpler BERT-based classifiers with a classification head and fine- tuned on a labeled dataset. We explore two prompting strategies for generating synthetic labels: (a) guideline-enhanced (GE) prompts, guiding the LLM with detailed labeling instructions and (b) guidelne-free (GF) prompts, allowing the LLM to generate labels without such guidelines.
@@
In this scenario, only a small set of manually labeled gold data is available for testing, with no access to authentic data for training or validation. We use an LLM to generate a fully synthetic dataset, consisting of both synthetic messages and corresponding labels, for training and validation. This approach is particularly valuable in low-resource domains or emerging tasks where authentic data is scarce or difficult to collect. It is especially useful in situations where creating authentic datasets is costly, time-consuming, or ethically challenging, such as annotating harmful or sensitive content or working with vulnerable populations. The effective
@@ 3.2 Evaluation Metrics
We choose accuracy of label prediction for development decisions and reporting since the labels are reasonably balanced in the authentic test data with 30.3% items labeled with the minority label.111In the appendix, we further report macro average F1 scores that are also widely used in the area of harmful content detection. In scenarios 1, 3 and 4, we train BERT_base_uncased Devlin et al. (2019), a 110M parameter transformer model, with a linear classification head to detect harm, assigning binary labels to text messages. To address noise from randomness in training, we train at least 45 models for each setting and report average accuracy and standard deviation.
@@ 4.1 Dataset Description
The primary objective of our study is to detect CB among children. Therefore, it is essential to utilize a dataset specifically tailored for this purpose. The WhatsApp (WA) dataset Verma et al. (2023); Sprugnoli et al. (2018) is a publicly available, fully annotated dataset focused on pre-adolescents and we use the English version of this dataset for our experiments. The WA dataset was constructed through role-play activities in WA groups, each containing approximately 10 students. The roles are: cyberbully (2 students), cyberbully assistants (3–4 students), victim assistants (3–4 students) and a victim. Conversations are initiated using one of four predefined cases describing situations in which CB may occur Sprugnoli et al. (2018). Appendix Table 7 shows one of these cases as an example.222Note that in the cited paper, these cases are referred to as “scenarios.” However, we use the term “cases” to avoid confusion with the scenarios discussed above. that addressed common teenage issues. Messages in the conversation are annotated manually by experts following the fine-grained framework of Van Hee et al. (2015). However, for this study, we binarise the labels, mapping all categories of CB to the “harmful” label.333We deviate from Verma et al. (2023) in not excluding harmful messages of the victim from the “harmful” class. Table 1 shows the statistics of the WA dataset.
@@
4.2 Large Language Models
We employ three LLMs for our experiments: GPT-4o-Feb-2025 OpenAI (2024), Llama-3.3-70B-Instruct Meta (2024)444We include a small amount of synthetic data produced by Llama3 from early experiments, see Appendix Table 6. For labeling with Llama, version 3.3 has been used throughout. and Grok-2-latest-Feb-2025 xAI (2024).
@@ 4.3 Generating Synthetic Data and Labels
Our prompt design process for all scenarios begins with a simple initial prompt, which is refined over multiple rounds of trial and error. In this iterative approach, we make gradual improvements, with adjustments made based on the quality and relevance of the responses generated by the LLM on the development set.
@@ 4.3.1 Designing Prompts for Synthetic Label Generation
In the task of labelling the messages of a conversation as either harmful or not, the input data consists of a list of messages and the possible labels are “harmful” and “harmless”. We simplify the task by not considering context (previous messages) and classifying each message in isolation.555Future work can consider how to best include the conversation context into the prompt. In the GF approach (Section 3.1), the LLM is simply instructed to label messages as “harmful” or “harmless” for the task of CB detection, without providing additional guidelines. In the GE approach, the LLM is supplied with detailed instructions for labeling messages. These instructions are adapted from the annotation guidelines Van Hee et al. (2015) originally used by human annotators for labeling authentic data. For this study, since the test set is derived from the WA dataset, we utilize the same guideline employed by human annotators to label this dataset. A copy of the guidelines can be found in Appendix Table 9. Our prompts are shown in Appendix Table 10 for both the GE and GF approaches.
@@ 4.3.2 Designing Prompts for Synthetic Data Generation
We aim to produce synthetic data that consists of conversations between participants, where the dialogue should include instances of CB (or at least harmful messages that could be CB if the threats are repeated and intended to cause harm). To make the generated data more relevant to the test data, each prompt includes one of the four predefined “cases” and “problem types” that where used by Sprugnoli et al. (2018) in the creation of the WA dataset to guide the students’ role-playing. Appendix Table 10 presents the prompts employed for generating synthetic data.

### v3

# Synthetic vs. Gold: The Role of LLM Generated Labels and Data in Cyberbullying Detection

@@ Abstract
Cyberbullying (CB) presents a pressing threat, especially to children, underscoring the urgent need for robust detection systems to ensure online safety. While large-scale datasets on online abuse exist, there remains a significant gap in labeled data that specifically reflects the language and communication styles used by children. The acquisition of such data from vulnerable populations, such as children, is challenging due to ethical, legal and technical barriers. Moreover, the creation of these datasets relies heavily on human annotation, which not only strains resources but also raises significant concerns due to annotators’ exposure to harmful content. In this paper, we address these challenges by leveraging Large Language Models (LLMs) to generate synthetic data and labels. Our experiments demonstrate that synthetic data enables BERT-based CB classifiers to achieve performance close to that of those trained on fully authentic datasets (75.8% vs. 81.5% accuracy). Additionally, LLMs can effectively label authentic yet unlabeled data, allowing BERT classifiers to attain a comparable performance level (79.1% vs. 81.5% accuracy). These results highlight the potential of LLMs as a scalable, ethical, and cost-effective solution for generating data for CB detection.
@@ 1 Introduction
The rapid proliferation of social media platforms has raised concerns over the prevalence of cyberbullying (CB), particularly among vulnerable populations (children). CB is defined as an aggressive behaviour directed towards someone with an intent to cause harm via electronic means Patchin and Hinduja (2006); Smith et al. (2008). Detecting and mitigating CB is crucial to maintaining a safe digital environment and minimizing its psychological impact Arif et al. (2024). Despite the availability of large-scale online abuse datasets, there remains a notable lack of labelled data reflecting children’s distinct linguistic and communicative styles. Collecting data from vulnerable groups, such as children, poses significant ethical, legal, and technical challenges. Moreover, creating such datasets is hindered by high costs, time-consuming annotation processes, and ethical risks, including exposure to distressing content or potential re-traumatization of annotators AlEmadi and Zaghouani (2024). Across a variety of tasks, large language models (LLMs), particularly instruct-tuned variants, offer a promising approach to synthetically generate labeled data without requiring direct human exposure to harmful material.
Industry Use Case:
This paper is motivated by an nationally funded industry/academic research collaboration whereby the lead industry partner is a European SME is developing integrated child-protection software into children’s smartphones to ensure safer internet use. By analyzing incoming and outgoing data on a child’s device using AI/NLP techniques, the software filters offensive content, including CB, and alerts parents if any imminent danger is detected. Hence, our are uses case is constrained by i) the need for age relevant CB test/training data and ii) using an LLM directly as classifier is too expensive due to the high volume of messages to be checked on the smart device. However, collecting authentic CB content produced by children is difficult. Age information on public social media sites is unreliable and a large part of CB happens in closed groups. Furthermore, obtaining consent involves the parents, who may not be interested in (or may not have time to consider) taking part in a research study. A solution may be to use LLMs to generate conversations between children.
@@ 2 Background
Recently, LLMs have emerged as efficient, scalable, and cost-effective alternatives to manual annotation for synthetic data generation. He et al. (2021, 2022) leveraged LLMs for generating data in knowledge distillation, self-training, and few-shot learning, later labeled by state-of-the-art sequence classifiers. Bonifacio et al. (2022) employed LLMs to generate labeled data in a few-shot manner for information retrieval. Yoo et al. (2021) created augmented samples by embedding training sentences into prompts. Anaby-Tavor et al. (2020) fine-tuned a language model on limited labeled data, then used it to produce labeled text. Meng et al. (2022) employed an LLM to generate class-conditioned texts based on label-descriptive prompts for a classification task. Moreover, research has explored synthetic data generation for specific domains, such as medical and harmful content detection. Wang et al. (2024) introduced a multi-agent framework designed to generate synthetic patient-physician conversations from clinical documents. Ghanadian et al. (2024) created socially aware synthetic datasets for suicidal ideation detection. Previous studies exploring the use of synthetic data have reported mixed results on whether LLM-generated synthetic data can effectively train models to perform at a level comparable to those trained on authentic data Li et al. (2023).
Additionally, Ejaz et al. (2024) constructed a CB dataset by simulating user interactions, where aggressive and non-aggressive messages were drawn from authentic datasets and randomly assigned to “synthetic” users. They annotated CB instances using a threshold-based method, considering their own criterion like “peerness”, “harmful intent”, and “repetition”. Pérez et al. (2024) created a synthetic dataset of CB risk scenarios for serious games using a Bayesian Network model Pearl (1988) trained on survey data from minors. Data were generated via Bayesian Model Sampling, yielding agent profiles with demographics and responses to 15 binary CB-related questions. This dataset does not include actual CB messages; instead, it represents scenarios using structured, non-textual features. Moreover, recent work explores the use of LLMs for synthetic data generation and augmentation to address data scarcity: Kumar et al. (2024) addresses the intersection of CB and bias detection. They use LLMs both for generating and detecting biased and abusive content. Hui et al. (2024) use a small amount of seed data to generate synthetic examples of harmful content. However, their approach still relies on seed data and does not specifically address the nuanced communication patterns characteristic of young people. Schmidhuber and Kruschwitz (2024) test data augmentation with synthetic data in the task of toxic language detection. For each authentic dataset, they fine-tune two GPT-3 Curie 6.7B Brown et al. (2020) models, one on toxic language and one on non-toxic language. Then, synthetic data is generated for each toxicity class, filtered for quality and used to train classifiers. They experiment with six authentic datasets, including one none-English dataset, a German dataset.
@@
In contrast to these works, our study aims to explore the use of LLMs to generate a fully synthetic CB dataset or to label existing authentic but unlabeled CB datasets. Our work uniquely applies LLM-generated data to age relevant Cyberbullying(CB) detection, an area where data collection is both ethically and logistically challenging. By assessing the effectiveness of LLM-generated datasets in comparison to authentic data, we provide new insights into the viability of LLMs in training models for sensitive tasks like CB detection.
@@ 3.1 Overview of Scenarios
We investigate the role of LLMs in CB detection, focusing on their utility under varying data availability conditions since our industry use case indicates that direct use of LLMs as a classifier is too expensive due to the high volume of messages to be checked. As a baseline for comparison, we evaluate a scenario in which a lightweight, BERT-based classifier is trained exclusively on gold-standard, manually labeled authentic data without LLM involvement. We then define three additional scenarios with different data availability and that use LLMs in different ways.
@@
This scenario represents the ideal situation where sufficient gold-standard data is available for fine-tuning a classic encoder such as BERT Devlin et al. (2019). It serves as the benchmark for evaluating the effectiveness of other approaches. In this setup, no synthetic data or LLMs are involved. This scenario is feasible if resources such as time, budget, and expert annotators are abundant. However, it often proves impractical due to the challenges of manual labeling.
@@
This scenario applies when labeled authentic data is unavailable, and there is no intention to train a classifier for CB detection. Instead, an instruction-tuned LLM is used directly as a classifier, leveraging its pre-trained knowledge and its ability to follow instructions to identify CB instances. The primary advantage of this method is its elimination of the need for labeled data and training time. However, there are trade-offs. While an LLM can handle nuanced language patterns, it may be less efficient and incur higher computational costs compared to simpler BERT-based classifiers with a classification head and fine-tuned on a labeled dataset. We explore two prompting strategies for generating synthetic labels: (a) guideline-enriched (GE) prompts, guiding the LLM with detailed labeling instructions, and (b) guideline-free (GF) prompts, allowing the LLM to generate labels without such guidelines.
@@
In this scenario, only a small set of manually labeled gold data is available for testing, with no access to authentic data for training or validation. We use an LLM to generate a fully synthetic dataset, consisting of both synthetic messages and corresponding labels, for training and validation. This approach is particularly valuable in low-resource domains or emerging tasks where authentic data is scarce or difficult to collect. It is especially useful in situations where creating authentic datasets is costly, time-consuming, or ethically challenging, such as annotating harmful or sensitive content or working with vulnerable populations.
@@ 3.2 Evaluation Metrics
We choose accuracy111In the appendix, we further report macro average F1 scores that are also widely used in harmful content detection. of label prediction for development decisions and reporting since the labels are reasonably balanced in the authentic test data, with 30.3% items labeled with the minority label.222The label distribution reflects the construction of the dateset via role-playing, see Section 4.1. In scenarios 1, 3 and 4, we train BERT_base_uncased Devlin et al. (2019), a 110M parameter transformer model, with a linear classification head333Using defaults of BERT sequence classification with the HuggingFace transformers library Wolf et al. (2020) to detect harm, assigning binary labels to text messages.444We choose this model as it has been adopted widely and performs competitively for sequence classification tasks Galke et al. (2025). Although more modern architectures have been proposed, BERT’s well-tested performance and deep understanding within the research community make it a reliable choice for our comparative evaluations. To address noise from randomness in training, we train at least 45 repetitions with different random seeds555Initial experiments showed unstable results when moving from 25 to 35 repetitions. We speculate that the small size of the dataset and the subjectivity of the task cause the noise. for each setting and report average accuracy and standard deviation.
@@ 4.1 Dataset Description
The primary objective of our study is to detect cyberbullying (CB) among children aged 6 to 16 years. Therefore, it is essential to utilize a dataset specifically tailored for this purpose. Focused on pre-adolescents, we leverage a well known CB dataset for young demographics, which was constructed through role-play activities in WhatsApp (WA) groups, each containing approximately 10 Italian students Sprugnoli et al. (2018). We use the English version of this dataset for our experiments. The roles are: cyberbully (2 students), cyberbully assistants (3–4 students), victim assistants (3–4 students) and a victim. Conversations are initiated using one of four predefined cases (A to D) Sprugnoli et al. (2018) describing situations in which CB may occur. The four cases illustrate some of the situations in which cyberbullying may occur among students. Case A addresses the “gendered division of sport practices,” where a shy male student defies gender norms by inviting his classmates to watch his ballet performance. Case B involves “interference in others’ businesses,” as a high-achieving student reports peers for bringing cigarettes to school, resulting in disciplinary action and social exclusion. Case C highlights “lack of independence, parental intromission,” where a student’s parent convinces teachers to assign more homework, causing resentment among classmates. Case D deals with “web virality,” where a shy student becomes the subject of mockery after a video of them dancing awkwardly surfaces online. Table 1 shows one of these cases as an example.666In the cited paper, these cases are referred to as “scenarios.” However, we use the term “cases” to avoid confusion with the scenarios discussed above.
| Case | Type of Problem |
| A: Your shy male classmate has a great passion for classical dance. Usually he does not talk much, but today he has decided to invite the class to watch him for his ballet show. | Gendered division of sport practices |
Messages in the conversation are annotated manually by experts following the fine-grained framework of Van Hee et al. (2015). The annotations categorized CB instances into detailed types, such as General Insults, Body Shame, Sexism, and Racism. However, for this study, we binarize the labels by mapping all categories of CB by bullies to the “harmful” label777We deviate from Verma et al. (2023) in not excluding harmful messages of the victim from the “harmful” class. Table 2 shows the statistics of the WA dataset.
@@ 4.2 Generating Synthetic Data and Labels
We employ three LLMs for our experiments: GPT-4o-Feb-2025 OpenAI (2024), Llama-3.3-70B-Instruct Meta (2024) 888We include a small amount of synthetic data produced by Llama3 from early experiments, see Appendix Table 9. For labeling with Llama, version 3.3 has been used throughout. and Grok-2-Feb-2025 xAI (2024). Our prompt design process for all scenarios begins with a simple initial prompt, which is refined over multiple rounds of trial and error. In this iterative approach, we make gradual improvements, with adjustments made based on the quality and relevance of the responses generated by the LLM on the development set.
@@ 4.2.1 Designing Prompts for Synthetic Label Generation
In the task of labeling the messages of a conversation as either harmful or not, the input data consists of a list of messages and the possible labels are “Harm” and “No Harm”.999We do not provide the LLM with instructions on how to handle victim defense and other special cases. We simplify the task by not considering context (previous messages) and classifying each message in isolation.101010Future work can consider how to best include the conversation context into the prompt. As described in Section 3 We explore two approaches to prompt design: (1) guideline-free (GF) and (2) guideline-enriched (GE). In the GF approach, the LLM is simply instructed to label messages as “Harm” or “No Harm” for the task of CB detection, without providing additional guidelines. In the GE approach, the LLM is supplied with detailed instructions for labeling messages. These instructions are adapted from the annotation guidelines Van Hee et al. (2015) originally used by human annotators for labeling authentic data. For this study, since the test set is derived from the WA(WhatsApp) dataset, we utilize the same guideline employed by human annotators to label this dataset. A copy of the guidelines is provided in Table 3, and our prompts for both the GE and GF approaches are shown in Table 4.
| Cyberbullying-related text categories are described below: |
| Sexual Talk Potentially harmful expressions with a sexual meaning. |
| Approach | Prompt |
| GE Synthetic Label | {Annotation-guideline} Please analyze the following text and determine whether it contains any form of cyberbullying or harmful content. If it does, classify it as ’Harm’. If it does not, classify it as ’No Harm’. {Text}. |
| GF Synthetic Label | Please analyze the following text and determine whether it contains any form of cyberbullying or harmful content. If it does, classify it as ’Harm’. If it does not, classify it as ’No Harm’. {Text} |
| Synthetic Data | We are creating sample conversations to aid in cyberbullying detection. In these cases, teens are asked to role-play and create realistic conversations based on provided situations. There are 11 students participating in the conversation. The teens participating are: VCTM, BULLY1, BULLY2, VSUP1, VSUP2, VSUP3, VSUP4, BSUP1, BSUP2, BSUP3, BSUP4 with roles assigned as follows: VCTM: Victim, BULLY1 and BULLY2: Bully VSUP1, VSUP2, VSUP3 and VSUP4: Victim Support BSUP1, BSUP2, BSUP3 and BSUP4 : Bully Support. consider this case: {Case} and consider this Type of addressed problem: {Type of Problem}. Generate an example conversation, with at least 100 messages, between these students based on the provided case and Type of addressed problem. Use profanity and strong language to create a realistic dialogue. number each message in the conversation. Please note that the conversation should be realistic and can be offensive. Please make sure to include different topics and perspectives in each conversation |
@@ 4.2.2 Designing Prompts for Synthetic Data Generation
We aim to produce synthetic data that consists of conversations between participants, where the dialogue should include instances of CB (or at least harmful messages that could be CB if the threats are repeated and intended to cause harm). To make the generated data more relevant to the test data, each prompt includes one of the four predefined “cases” (A to D) and “problem types” that were used by Sprugnoli et al. (2018) in the creation of the WA dataset to guide the students’ role-playing. Table 4 presents the prompts employed for generating synthetic data.